OpenStudy (anonymous):

Solve the equation algebraically. e^(.05t)=3

OpenStudy (anonymous):

t = (ln(3))/.5

OpenStudy (anonymous):

ln(e^.05t)=ln(3) 0.05t=ln(3) t=ln(3)/(0.05)